# Variable Elliptic Structures on the Plane: Transport Dynamics, Rigidity, and Function Theory

### v4

@@ 8.1 The Rigid Variable Laplace Operator
In the present framework, once rigidity is imposed (Chapter 2), $\partial_{\bar{z}}$ becomes a derivation and the Cauchy–Pompeiu formula (Chapter 5) becomes available. It is then natural—and unavoidable—to study the second–order operator obtained by composing $\partial_{\bar{z}}$ with its conjugate partner $\partial_{z}$.
@@ 9.1.2 Canonical principal part
By elementary substitutions preserving ellipticity, system (9.1) can be written in the form
@@
where $a,b,c,d,f,g$ are real-valued functions on $\Omega$.
@@
$W:=U+iV.$
9.1.3 Intrinsic decomposition of $\bar{\partial}W$
Define the variable–structure Cauchy–Riemann operator
@@ 9.1.3 Intrinsic decomposition of $\bar{\partial}W$
Substituting (9.3) into (9.4) gives
$2\bar{\partial}W-VG=(f+ig)-(a+ic)U-(b+id)V.$ (9.5)
@@ 9.1.4 Elliptic conjugation and Vekua form
Since $\Delta>0$, the element $2i+\beta$ is invertible in $A_{z}$ and in fact
$(2i+\beta)^{-1}=\frac{\beta-2i}{\beta^{2}+4\alpha}.$
@@
Substituting (9.7) into the right–hand side of (9.5) expresses it as an $A_{z}$–linear combination of $W$ and $\widehat{W}$. Concretely, there exist unique algebra–valued coefficients $A_{0},B_{0}$ and a unique datum $F_{0}$ such that
$(f+ig)-(a+ic)U-(b+id)V\;=\;2F_{0}-2A_{0}\,W-2B_{0}\,\widehat{W}$. (9.8)
(Existence and uniqueness follow because (9.7) is an invertible linear change of variables between $(U,V)$ and (W,W^)($\widehat{W}$).)
Define
$F:=F_{0}=\tfrac{1}{2}(f+ig),$
and absorb the obstruction term $VG$ from (9.5) using $V=(W-\widehat{W})/(2i+\beta)$:
$VG=\frac{W-\widehat{W}}{2i+\beta}\,G.$
Thus (9.5) may be rewritten uniquely in the form
$2\bar{\partial}W=2F-2A\,W-2B\,\widehat{W}$, (9.9)
where
$A\;:=\;A_{0}-\frac{1}{2}\,\frac{G}{2i+\beta},\qquad B\;:=\;B_{0}+\frac{1}{2}\,\frac{G}{2i+\beta}.$ (9.10)
@@ 9.1.5 Remarks
  • Equation (9.11) is intrinsic to the elliptic structure $(\alpha,\beta)$ and requires no auxiliary reduction.
  • When the structure is rigid ($G\equiv 0$), the obstruction contribution in (9.10) vanishes and the lower–order coefficients reduce to the rigid forms determined by (9.8).
  • The obstruction $G$ measures precisely the deviation from rigid Cauchy–Riemann calculus and is absorbed into the lower–order terms via the explicit splitting in (9.10).
@@ 9.2.1 Rigid and non–rigid structures
Equivalently, rigidity means that the variable–structure Cauchy–Riemann operator $\bar{\partial}=\tfrac{1}{2}(\partial_{x}+i\,\partial_{y})$ satisfies the Leibniz rule on algebra–valued functions.
@@ 9.3 Rigid Similarity Principle (Sharp Case $F=0$, $B=0$)
$G\equiv 0,$
so that the variable–structure Cauchy–Riemann operator
$\bar{\partial}:=\tfrac{1}{2}(\partial_{x}+i\partial_{y})$
satisfies the Leibniz rule on algebra–valued functions.
@@ 9.5 Initial Value Problems for Rigid Vekua Systems
Throughout this section we assume that the elliptic structure is rigid,
$G\equiv 0,$
so that the variable–structure Cauchy–Riemann operator
$\bar{\partial}=\tfrac{1}{2}(\partial_{x}+i\partial_{y})$
satisfies the Leibniz rule and the full rigid calculus applies.
@@ 12.1 Purpose of This Chapter
  • the normalized obstruction $\rho_{T}$ that controls rigidization convergence (Chapter 10) is the norm of a Poincaré residual weighted by the hyperbolic metric density (Proposition 12.8).
These observations connect the transport theory of Chapters 23 to the classical hyperbolic geometry of the space of conformal structures.
@@ 12.4.1 Dilatation of the structure map
12.5 The Poincaré Residual and the Normalized Obstruction
@@ 12.5.1 Definition of the residual
Definition 12.7 (Poincaré residual).
The Poincaré residual of a variable elliptic structure is
@@ 12.5.2 Connection to the normalized obstruction
where $b_{\min}=\min_{\bar{\Omega}}\operatorname{Im}\lambda$ is the spectral floor. We now express $\rho_{T}$ in terms of the Poincaré residual.
@@ 12.5.3 The differential of a rigid structure
The Poincaré residual admits a differential-geometric reformulation. Write $d\mu=\mu_{z}\,dz+\mu_{\bar{z}}\,d\bar{z}$. The rigidity condition $\mu_{\bar{z}}=\mu\,\mu_{z}$ is equivalent to
@@ 12.7.1 Teichmüller theory
The disk $\mathbb{D}$ parameterizes conformal structures on the plane, and the Poincaré metric on $\mathbb{D}$ induces the Teichmüller metric on deformation spaces. The rigidity condition (12.7) characterizes the variable structures $\mu:\Omega\to\mathbb{D}$ that are “self-consistent” in the sense that their variation across $\Omega$ is compatible with the conformal geometry they define. The Poincaré residual $\mathcal{R}$ measures the failure of this compatibility, weighted by the Teichmüller distance to the boundary of moduli space.
@@ 12.8 The Independence Theorem, Revisited
Let $\Omega\subset\mathbb{R}^{2}$ be a domain. The Poincaré residual $\|\mathcal{R}(\mu)\|_{\infty}$ and the supremum $\|\mu\|_{\infty}$ are independently prescribable: for any $R_{0}\geq 0$ and $0\leq\mu_{0}<1$, there exists a $C^{\infty}$ map $\mu:\Omega\to\mathbb{D}$ with $\|\mathcal{R}\|_{\infty}$ arbitrarily close to $R_{0}$ and $\|\mu\|_{\infty}$ arbitrarily close to $\mu_{0}$.
@@
The residual $\mathcal{R}=\mu_{\bar{z}}-\mu\,\mu_{z}$ depends on the first derivatives of $\mu$, while $|\mu|$ depends on the pointwise values. Since these are different orders of the jet, they are independently prescribable: one can construct maps with prescribed pointwise values and prescribed first derivatives (subject only to smoothness). The explicit Burgers solutions of Chapter 4 provide $\mathcal{R}=0$ with arbitrary $\|\mu\|_{\infty}$, and perturbations of $\mu=0$ provide arbitrary $\|\mathcal{R}\|_{\infty}$ with $\|\mu\|_{\infty}\approx 0$. ∎
@@ 12.9 Comparison with the Spectral Formulation
Equation type real Burgers complex Beltrami–Burgers
@@ 12.10 Discussion and Open Questions
The normalized obstruction $\rho_{T}$, which controls the convergence of the rigidization scheme (Chapter 10, Theorem 10.15), turns out to measure the Poincaré residual weighted by the hyperbolic metric density. This was not designed; it emerged from the algebra of the Möbius transform relating $\lambda$ and $\mu$. The hyperbolic weight $(1-|\mu|^{2})^{-2}$ was present in the original theory as the normalization $b_{\min}^{-2}$, where $b_{\min}=\min\operatorname{Im}\lambda$ is the spectral floor. The Poincaré metric was hiding in plain sight.
@@
[…]
@@ 13.1 Purpose of This Chapter
    The transport field $R(\mu)=\mu_{\bar{z}}-\mu\,\mu_{z}$, the Poincaré residual of Definition 12.6, which measures how the structure moves through the disk.
@@ 13.2 Two Invariants at Different Orders
The Poincaré residual $R(\mu)=\mu_{\bar{z}}-\mu\,\mu_{z}$ (Definition 12.6; recalled here for the reader’s convenience) depends on the first derivatives of $\mu$. Its Hölder norm $\|R(\mu)\|_{C^{0,\alpha}(\overline{\Omega})}$ measures how the structure moves through the Poincaré disk.
The Poincaré residual is the invariant that the rigidizer sees. Theorem 10.15 converges when $\|R\|_{C^{0,\alpha}}$ is small relative to the Poincaré metric density at the structure’s position; its convergence is blind to $\|\mu\|_{C^{0}}$ except through this metric weighting.

### v5

@@ 8.1 The Rigid Variable Laplace Operator
In the present framework, once rigidity is imposed (Chapter 2), $\partial_{\bar{z}}$ becomes a derivation and the Cauchy–Pompeiu formula (Chapter 6) becomes available. It is then natural—and unavoidable—to study the second–order operator obtained by composing $\partial_{\bar{z}}$ with its conjugate partner $\partial_{z}$.
@@ 9.1.1 Elliptic system and induced structure
9.1.2 The structure polynomial is determined by the principal symbol
The coefficients $\alpha,\beta$ defined in (9.2) are not an artifact of the substitution used to reach canonical form. They are determined, uniquely and intrinsically, by the principal symbol of the elliptic system.
Proposition 9.1.
Let $\sigma(z;\xi,\eta)$ denote the principal symbol of (9.1) at the point $z\in\Omega$ and covector $(\xi,\eta)\in\mathbb{R}^{2}$. Then
$\det\sigma(z;\xi,\eta)=a_{11}\bigl(\xi^{2}-\beta\,\xi\eta+\alpha\,\eta^{2}\bigr)=a_{11}\,N(\xi+i\eta),$ (9.3)
where $N$ is the algebra norm of $A_{z}$, i.e. the norm form determined by the structure polynomial $X^{2}+\beta X+\alpha$.
Proof.
The principal symbol of (9.1) is the matrix
$\sigma(z;\xi,\eta)=\begin{pmatrix}a_{11}\xi+a_{12}\eta&-\eta\\[4.0pt] a_{21}\xi+a_{22}\eta&\xi\end{pmatrix}$.
Its determinant is
$\det\sigma=(a_{11}\xi+a_{12}\eta)\,\xi+(a_{21}\xi+a_{22}\eta)\,\eta=a_{11}\xi^{2}+(a_{12}+a_{21})\,\xi\eta+a_{22}\,\eta^{2}.$
Dividing by $a_{11}>0$ and substituting (9.2):
$\frac{\det\sigma}{a_{11}}=\xi^{2}+\frac{a_{12}+a_{21}}{a_{11}}\,\xi\eta+\frac{a_{22}}{a_{11}}\,\eta^{2}=\xi^{2}-\beta\,\xi\eta+\alpha\,\eta^{2}.$ (9.4)
On the other hand, in the algebra $A_{z}$ with generator satisfying $i^{2}+\beta\,i+\alpha=0$, the norm of an element $w=a+ib$ is
$N(a+ib)=w\cdot\widehat{w}=(a+ib)(a+\widehat{i}\,b)=a^{2}-\beta\,ab+\alpha\,b^{2}.$
Evaluating at $w=\xi+i\eta$ gives
$N(\xi+i\eta)=\xi^{2}-\beta\,\xi\eta+\alpha\,\eta^{2},$
which coincides with (9.4). ∎
This identification has three immediate consequences.
  1. 1.
    Ellipticity is equivalent to ellipticity of the algebra.  The system (9.1) is elliptic (i.e. $\det\sigma\neq 0$ for all $(\xi,\eta)\neq(0,0)$) if and only if the norm form $N(\xi+i\eta)$ is positive definite, which is precisely the condition $\Delta=4\alpha-\beta^{2}>0$.
  2. 2.
    Uniqueness.  For a given elliptic system, the structure polynomial $X^{2}+\beta X+\alpha$ is the unique monic quadratic whose associated norm form equals the normalized symbol determinant. The coefficients $\alpha,\beta$ are therefore invariants of the principal part, not of the reduction procedure.
  3. 3.
    The substitution realizes the symbol.  The explicit substitution in Lemma 9.4 does not choose $\alpha$ and $\beta$; it produces a system whose principal part is written in the algebra already determined by the symbol.
Corollary 9.2.
The structure polynomial $X^{2}+\beta X+\alpha$ is invariant under elliptic–compatible row operations and invertible substitutions on the target variables.
Proof.
Let $R=R(z)$ be an invertible $2\times 2$ matrix (a row operation on the system) and $M=M(z)$ an invertible substitution on the target variables $(u,v)\mapsto M(u,v)$. Under these transformations the principal symbol changes as
$\sigma\;\longmapsto\;R\,\sigma\,M^{-1},$
so
$\det\sigma\;\longmapsto\;\frac{\det R}{\det M}\,\det\sigma.$
The prefactor $\det R/\det M$ is a nonzero scalar depending only on $z$, not on the covector $(\xi,\eta)$. Since $\alpha$ and $\beta$ are obtained by normalizing $\det\sigma$ to monic form in $\xi$ (i.e. dividing by the coefficient of $\xi^{2}$), the scalar cancels identically:
$\frac{\det(R\,\sigma\,M^{-1})}{a_{11}\,\det R/\det M}=\frac{\det\sigma}{a_{11}}=\xi^{2}-\beta\,\xi\eta+\alpha\,\eta^{2}.$
Therefore $\alpha$ and $\beta$ depend only on the equivalence class of the system, not on the particular representative. ∎
Remark 9.3.
In particular, the substitution $U=a_{22}\,u$, $V=v-a_{12}\,u$ used in Lemma 9.4 is one choice of representative within this equivalence class. Any other elliptic–compatible reduction would produce a system governed by the same structure polynomial.
@@ 9.1.3 Canonical principal part
Lemma 9.4.
Suppose the principal coefficients $a_{ij}$ in (9.1) are of class $C^{1}$ with $a_{11},a_{22}>0$. Then the substitution
$U=a_{22}\,u,\qquad V=v-a_{12}\,u,$
followed by a single row operation and a scaling, reduces (9.1) to the canonical form
@@
where $a,b,c,d,f,g$ are real–valued $C^{0}$ functions determined by the original coefficients and their first derivatives.
Proof.
Under the substitution $U=a_{22}\,u$, $V=v-a_{12}\,u$ (with $a_{22}>0$), the principal parts of (9.1) become
$\begin{cases}-V_{y}+\dfrac{a_{11}}{a_{22}}\,U_{x}+a_{\ast}U+b_{\ast}V=f_{1},\\[8.0pt] \ \ V_{x}+\dfrac{a_{21}+a_{12}}{a_{22}}\,U_{x}+U_{y}+c_{\ast}U+d_{\ast}V=f_{2},\end{cases}$ (9.6)
where $a_{\ast},b_{\ast},c_{\ast},d_{\ast}$ absorb all contributions from the derivatives of the variable coefficients $a_{ij}$ and the original lower–order terms. With the structure coefficients (9.2) we have
$\frac{a_{11}}{a_{22}}=\frac{1}{\alpha},\qquad\frac{a_{21}+a_{12}}{a_{22}}=-\frac{\beta}{\alpha},$
so (9.6) reads
$\begin{cases}-V_{y}+\dfrac{1}{\alpha}\,U_{x}+a_{\ast}U+b_{\ast}V=f_{1},\\[8.0pt] \ \ V_{x}-\dfrac{\beta}{\alpha}\,U_{x}+U_{y}+c_{\ast}U+d_{\ast}V=f_{2}.\end{cases}$ (9.7)
Row operation.  Multiply the first equation of (9.7) by $\beta$ and add to the second. The $U_{x}$ coupling in the second line cancels:
$\frac{\beta}{\alpha}\,U_{x}+\frac{\beta}{\alpha}\,U_{x}=0,$
yielding
$\begin{cases}-V_{y}+\dfrac{1}{\alpha}\,U_{x}+a_{\ast}U+b_{\ast}V=f_{1},\\[8.0pt] \ \ V_{x}-\beta V_{y}+U_{y}+c^{\prime}U+d^{\prime}V=g^{\prime}.\end{cases}$ (9.8)
Scaling.  Multiply the first equation of (9.8) by $\alpha>0$ to obtain the canonical form (9.5). ∎
Remark 9.5.
The structure coefficients $\alpha$ and $\beta$ are not imposed on the system; they are read off from it. No uniform ellipticity assumption is required, and no auxiliary Beltrami equation is solved.
9.1.4 Intrinsic decomposition and the obstruction
@@
$W:=U+iV,$
and define the variable–structure Cauchy–Riemann operator
@@
Substitution.  The canonical system (9.5) identifies the two real components of (9.9) directly:
$U_{x}-\alpha V_{y}$ $=f-aU-bV,$
$V_{x}+U_{y}-\beta V_{y}$ $=g-cU-dV.$
Substituting into (9.9) gives
$2\bar{\partial}W=(f-aU-bV)+(g-cU-dV)\,i+V\,G.$ (9.10)
Decomposing the obstruction.  Write $G$ in its real and imaginary components in the algebra $A_{z}$:
$G=G_{1}+G_{2}\,i,\qquad G_{1},\,G_{2}\;\text{real.}$
Since $V$ is real, $VG=VG_{1}+(VG_{2})\,i$. Adding the obstruction components to each line of (9.10):
$2\bar{\partial}W=\bigl(f-aU-(b-G_{1})V\bigr)+\bigl(g-cU-(d-G_{2})V\bigr)\,i.$ (9.11)
Define the modified lower–order coefficients
$\tilde{b}:=b-G_{1},\qquad\tilde{d}:=d-G_{2}.$
Then (9.11) reads
$2\bar{\partial}W=(f+ig)-(a+ic)\,U-(\tilde{b}+i\tilde{d})\,V.$ (9.12)
Remark 9.6.
When the structure is rigid ($G\equiv 0$), we have $G_{1}=G_{2}=0$ and $\tilde{b}=b$, $\tilde{d}=d$. In the general case, the obstruction modifies only the coefficients of $V$ and does so additively.
@@ 9.1.5 Elliptic conjugation and Vekua form
Since $\Delta>0$, the element $2i+\beta$ is invertible in $A_{z}$, with
$(2i+\beta)^{-1}=\frac{\beta-2i}{\beta^{2}+4\alpha}=\frac{\beta-2i}{\Delta}.$
@@
Substituting (9.14) into the right–hand side of (9.12) expresses it as an $A_{z}$–linear function of $W$ and $\widehat{W}$. Since (9.14) is an invertible linear change of variables between $(U,V)$ and (W,W^)($\widehat{W}$), there exist unique algebra–valued coefficients $A,B$ and a unique datum $F$ such that
$2\bar{\partial}W=2F-2A\,W-2B\,\widehat{W}$. (9.15)
@@
For reference, the coefficients are given explicitly by
$A$ $=\frac{1}{4}\!\left(a-\frac{\beta}{\alpha}\,\tilde{b}+\tilde{d}\;+\;\Bigl(c-\frac{1}{\alpha}\,\tilde{b}\Bigr)\,i\right),$ (9.17)
$B$ $=\frac{1}{4}\!\left(a+\frac{\beta}{\alpha}\,\tilde{b}-\tilde{d}\;+\;\Bigl(c+\frac{1}{\alpha}\,\tilde{b}\Bigr)\,i\right),$
$F$ $=\tfrac{1}{2}(f+ig),$
where $\tilde{b}=b-G_{1}$, $\tilde{d}=d-G_{2}$, and $G=G_{1}+G_{2}\,i$ is the intrinsic obstruction.
@@ 9.1.6 Remarks
  • Equation (9.16) is intrinsic to the elliptic structure $(\alpha,\beta)$ and requires no auxiliary Beltrami reduction.
  • The entire derivation is constructive: the substitution $(u,v)\mapsto(U,V)$ and the row operation are explicit, and the structure coefficients $(\alpha,\beta)$ emerge directly from the principal part of the system.
  • The obstruction $G=G_{1}+G_{2}\,i$ enters only by modifying the lower–order coefficients of $V$: it shifts $b\mapsto\tilde{b}=b-G_{1}$ and $d\mapsto\tilde{d}=d-G_{2}$. When the structure is rigid ($G\equiv 0$), the coefficients $A,B$ in (9.17) reduce to those of the classical Vekua theory with structure polynomial $X^{2}+\beta X+\alpha$.
@@ 9.3 Rigid Similarity Principle (Sharp Case $F=0$, $B=0$)
$G\equiv 0$
@@ 12.1 Purpose of This Chapter
  • the normalized obstruction $\rho_{T}$ that controls rigidization convergence (Chapter 10) is the norm of a Beltrami residual weighted by the hyperbolic metric density (Proposition 12.8).
These observations connect the transport theory of Chapters 24 to the classical hyperbolic geometry of the space of conformal structures.
@@ 12.4.1 Dilatation of the structure map
12.5 The Beltrami Residual and the Normalized Obstruction
@@ 12.5.1 Definition of the residual
Definition 12.7 (Beltrami residual).
The Beltrami residual of a variable elliptic structure is
@@ 12.5.2 Connection to the normalized obstruction
where $b_{\min}=\min_{\bar{\Omega}}\operatorname{Im}\lambda$ is the spectral floor. We now express $\rho_{T}$ in terms of the Beltrami residual.
@@ 12.5.3 The differential of a rigid structure
The Beltrami residual admits a differential-geometric reformulation. Write $d\mu=\mu_{z}\,dz+\mu_{\bar{z}}\,d\bar{z}$. The rigidity condition $\mu_{\bar{z}}=\mu\,\mu_{z}$ is equivalent to
@@ 12.7.1 Teichmüller theory
The disk $\mathbb{D}$ parameterizes conformal structures on the plane, and the Poincaré metric on $\mathbb{D}$ induces the Teichmüller metric on deformation spaces. The rigidity condition (12.7) characterizes the variable structures $\mu:\Omega\to\mathbb{D}$ that are “self-consistent” in the sense that their variation across $\Omega$ is compatible with the conformal geometry they define. The Beltrami residual $\mathcal{R}$ measures the failure of this compatibility, weighted by the Teichmüller distance to the boundary of moduli space.
@@ 12.8 The Independence Theorem, Revisited
Let $\Omega\subset\mathbb{R}^{2}$ be a domain. The Beltrami residual $\|\mathcal{R}(\mu)\|_{\infty}$ and the supremum $\|\mu\|_{\infty}$ are independently prescribable: for any $R_{0}\geq 0$ and $0\leq\mu_{0}<1$, there exists a $C^{\infty}$ map $\mu:\Omega\to\mathbb{D}$ with $\|\mathcal{R}\|_{\infty}$ arbitrarily close to $R_{0}$ and $\|\mu\|_{\infty}$ arbitrarily close to $\mu_{0}$.
@@
The residual $\mathcal{R}=\mu_{\bar{z}}-\mu\,\mu_{z}$ depends on the first derivatives of $\mu$, while $|\mu|$ depends on the pointwise values. Since these are different orders of the jet, they are independently prescribable: one can construct maps with prescribed pointwise values and prescribed first derivatives (subject only to smoothness). The explicit Burgers solutions of Chapter 5 provide $\mathcal{R}=0$ with arbitrary $\|\mu\|_{\infty}$, and perturbations of $\mu=0$ provide arbitrary $\|\mathcal{R}\|_{\infty}$ with $\|\mu\|_{\infty}\approx 0$. ∎
@@ 12.9 Comparison with the Spectral Formulation
Equation type complex Burgers complex Beltrami–Burgers
@@ 12.10 Discussion and Open Questions
The normalized obstruction $\rho_{T}$, which controls the convergence of the rigidization scheme (Chapter 10, Theorem 10.15), turns out to measure the Beltrami residual weighted by the hyperbolic metric density. This was not designed; it emerged from the algebra of the Möbius transform relating $\lambda$ and $\mu$. The hyperbolic weight $(1-|\mu|^{2})^{-2}$ was present in the original theory as the normalization $b_{\min}^{-2}$, where $b_{\min}=\min\operatorname{Im}\lambda$ is the spectral floor. The Poincaré metric was hiding in plain sight.
@@
$\text{Beltrami residual}\;\to\;\text{self-dilatation}\;\to\;\text{analytic closure}.$
@@ 13.1 Purpose of This Chapter
    The transport field $R(\mu)=\mu_{\bar{z}}-\mu\,\mu_{z}$, the Beltrami residual of Definition 12.6, which measures how the structure moves through the disk.
@@ 13.2 Two Invariants at Different Orders
The Beltrami residual $R(\mu)=\mu_{\bar{z}}-\mu\,\mu_{z}$ (Definition 12.6; recalled here for the reader’s convenience) depends on the first derivatives of $\mu$. Its Hölder norm $\|R(\mu)\|_{C^{0,\alpha}(\overline{\Omega})}$ measures how the structure moves through the Poincaré disk.
The Beltrami residual is the invariant that the rigidizer sees. Theorem 10.15 converges when $\|R\|_{C^{0,\alpha}}$ is small relative to the Poincaré metric density at the structure’s position; its convergence is blind to $\|\mu\|_{C^{0}}$ except through this metric weighting.